[Brahms-l] run update

From: flemming videbaek <videbaek@rcf.rhic.bnl.gov>
Date: Mon Feb 02 2004 - 15:11:36 EST
Just a brief update on the run status.

- The survey of soft-pt physics , in particular for protons and anti-protons, is complete, in so far the first
  round of 12,10,8,6, and 4 deg lower field settings is done. 
  The plan is to move the FS during the Wednesday access from the
  current 3 deg to make the 2.3 deg measurements. 
  The data taken so far should be reviewed for completenes in terms of settings
  problems etc.. so follow-up runs can be taken later.
- The 3 deg running at full field had just started, unfortunately a fuse for the D2 power- supply is blown, and no
  on-site replacement exists. It looks as if it at the earliest can come back some time tomorrow, which in ernest means
  physics running again following the Tuesday eve+night machine exp + 12 hours access i.e. late wednesday.
  The options are being reviewed by Dana and the CA-D engineer.
- Until then the program will just be high field running at 40 deg.

flemming

----------------------------------------------------------------
Flemming Videbaek
Physics Department
Brookhaven National Laboratory

e-mail: videbaek@bnl.gov
phone: 631-344-4106



_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Feb 2 15:14:41 2004

This archive was generated by hypermail 2.1.8 : Mon Feb 02 2004 - 15:15:06 EST